OpenStudy (anonymous):

Graph f(x)=2x^2-4, and find the inverse

OpenStudy (anonymous):

@peachpi @vera_ewing

OpenStudy (anonymous):

Did you graph it? and is there a domain restriction for the inverse?

OpenStudy (anonymous):

yes, i graphed it. What does that mean 'domain restriction'?

OpenStudy (anonymous):

is there a restriction on the x values? I ask because f(x) is a parabola and it isn't one-to-one which generally means it doesn't have an inverse without the restriction

OpenStudy (anonymous):

http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiIyeF4yLTQiLCJjb2xvciI6IiMwMDAwMDAifSx7InR5cGUiOjEwMDB9XQ-- This is the graph, so not that I can see

OpenStudy (anonymous):

You can still solve for an inverse using the usual process, but you can only do it for half the function at a time I think. f(x) = 2x² - 4 y = 2x² - 4 Switch x and y and solve for y x = 2y² - 4 x + 4 = 2y² ½x + 2 = y² \[y=\pm \sqrt{\frac{ 1 }{ 2 }x+2}\]

OpenStudy (anonymous):

okay:) I understand that :)

OpenStudy (anonymous):

How would you graph the inverse?

OpenStudy (anonymous):